# Think2SQL: Reinforce LLM Reasoning Capabilities for Text2SQL

### v1

@@ 4.3 Ablation on different training strategies
Table 2 presents an ablation study comparing different training strategies applied to Qwen2.5-Coder models of two sizes (3B and 7B parameters). Across both model sizes, we observe that applying SFT with reasoning traces consistently improves performance over the base model. Notably, the SFTNO-THINK variant, which lacks reasoning traces, shows a significant drop in performance compared to the full SFT model, particularly on the more challenging examples. For the 3B model, the SFT variant outperforms the SFTNO-THINK of 0.04 points on average, indicating that the reasoning traces are beneficial for the model’s performance.
The model trained only with RL further boosts performance, especially on the more challenging subsets. The combination of SFT and RL (SFT-RLEX) leads to robust and balanced improvements, although in some cases, pure RLEX slightly outperforms it.
For the larger Qwen2.5-Coder-7B model, we see a clear performance gain over its 3B counterpart across all difficulty levels, indicating that model scale remains a strong factor especially for RL when trained with GRPO. The best overall result is obtained with Qwen2.5-Coder-7B-RLEX, achieving the highest average score (0.552) and the strongest performance on Simple and Medium and Challenging examples.
Takeaway 2: RL vs SFT SFT improves over base models, especially for smaller LLMs. RL improves performance across all difficulty levels, particularly in more complex scenarios.
Table 3: Analysis robustness for different datasets.
@@
Qwen2.5-Coder-3B-RLEX $0.777$ $0.717$ 0.680
Qwen2.5-Coder-3B-SFT-RLEX 0.783 0.760 $0.658$
@@
Qwen2.5-Coder-7B-SFT-RLEX 0.824 0.780 $0.673$
@@ 5 Conclusions, limitations, and future work
This paper investigated the influence of reasoning capabilities on the performance of LLMs for the Text2SQL task. We evaluated different training strategies—Zero-Shot Learning (ZSL) with and without general-purpose reasoning, Supervised Fine-Tuning (SFT) with and without task-specific reasoning traces, Reinforcement Learning (RL) optimizing execution accuracy, and a combined SFT+RL approach—across multiple benchmark datasets.
Our findings answer three research questions. RQ1: while general-purpose reasoning in pretrained LLMs offers limited benefits for complex Text2SQL under ZSL, incorporating task-specific reasoning traces via SFT significantly improves performance, particularly for smaller models. RQ2: RL proved highly effective across all models and datasets, especially for queries demanding multi-hop reasoning - pure RL often yielded the best performance on challenging subsets. RQ3: the combined SFT+RL strategy demonstrated strong generalization across diverse datasets, suggesting it strikes an effective balance between learning general reasoning patterns (via SFT) and optimizing for task-specific correctness (via RL). Our Think2SQL-7B model, trained with RL, achieved performance on par with models exceeding 100 billion parameters on the BIRD dataset, showcasing the power of targeted reasoning reinforcement.
Despite these promising results, this study has limitations. We deliberately isolated the SQL generation process by providing the relevant database schema subset, excluding the challenge of automated schema linking, which is critical in real-world applications. Our RL reward function, while effective, primarily relied on binary execution accuracy, which can be sparse. Furthermore, the SFT training relied on synthetically generated reasoning traces, whose style and quality might influence outcomes. The core experiments focused on the Qwen-Coder-2.5 family, and findings might vary across different model architectures.
Future work should focus on exploring more sophisticated reward modeling for RL, potentially incorporating learned reward models or rewarding partial query correctness to further enhance training efficiency and performance. Evaluating these reasoning-focused training strategies on a wider range of LLM architectures and sizes, as well as on diverse, potentially proprietary or enterprise-specific datasets, would also be valuable. Finally, a deeper qualitative analysis of model failure cases could provide further insights into how different reasoning strategies succeed or falter on specific types of complex queries.

### v2

@@ 4.3 Ablation on different training strategies
Table 2 presents an ablation study comparing different training strategies applied to Qwen2.5-Coder models of two sizes (3B and 7B parameters). Across both model sizes, we observe that applying SFT with reasoning traces consistently improves performance over the base model. Notably, the SFTNT variant, which lacks reasoning traces, shows a significant drop in performance compared to the full SFT model, particularly on the more challenging examples. For the 3B model, the SFT variant outperforms the SFTNT of 0.04 points on average, indicating that the reasoning traces are beneficial for the model’s performance. Instead, for the 7B model, the SFTNT variant performs slightly worse than the base model suggesting possible overfitting to the training data. This highlights the importance of reasoning traces in guiding the model’s learning process.
The model trained exclusively with RL exhibits further performance gains, particularly on the more challenging subsets. Among the reward functions considered, $R_{\text{QATCH}}$ yields the best results, consistently outperforming $R_{\text{EX}}$ across both the 3B and 7B model sizes. These findings suggest that sparse rewards such as $R_{\text{EX}}$ are less effective in guiding the model’s learning process compared to denser, more informative reward signals.
For the larger Qwen2.5-Coder-7B model, we observe a clear performance gain over its 3B counterpart across all difficulty levels, confirming that model scale remains a significant factor, particularly when fine-tuned with RL via GRPO. The best overall results are achieved by Qwen2.5-Coder-7B-RLQATCH, which attains the highest average score (0.561) and the strongest performance on both Simple and Medium examples. While Qwen2.5-Coder-7B-RLEX slightly outperforms on Challenging examples, this advantage is marginal and must be interpreted with caution, given the limited size of the Challenging set (only 143 samples). Notably, the $R_{\text{EX}}$-trained model correctly solves only two additional examples, highlighting that $R_{\text{QATCH}}$ offers a more consistent and robust improvement across the broader evaluation spectrum.
Takeaway 2: RL vs SFT SFT improves over base models, especially for smaller LLMs. RL improves performance across all difficulty levels, particularly in more complex scenarios. Dense Rewards as $R_{\text{QATCH}}$ are more effective than sparse rewards like $R_{\text{EX}}$ for RL .
Table 3: Analysis robustness for different datasets. In bold are the best results for each model size. The RLEX and RLQATCH are the RL training strategies with $R_{\text{EX}}$ and $R_{\text{QATCH}}$, respectively. In parenthesis, we show the relative improvement over the base model.
@@
Qwen2.5-Coder-3B-RLEX $0.780$ $0.719$ $0.648$
Qwen2.5-Coder-3B-RLQATCH $0.777$ $0.717$ 0.680 (+7.2%)
Qwen2.5-Coder-3B-SFT-RLEX 0.783 (+5.8%) 0.760 (+12.8%) $0.658$
@@
Qwen2.5-Coder-7B-RLQATCH $0.822$ $0.769$ $\textbf{0.731}{\small\leavevmode\nobreak\ (+7.9\%)}$
Qwen2.5-Coder-7B-SFT-RLEX $\textbf{0.824}{\small\leavevmode\nobreak\ (+5.8\%)}$ 0.780 (+5.8%) $0.673$
@@ 5 Conclusions, limitations, and future work
This paper investigated the influence of reasoning capabilities on the performance of LLMs for the Text2SQL task. We evaluated different training strategies—Zero-Shot Learning (ZSL) with and without general-purpose reasoning, Supervised Fine-Tuning (SFT) with and without task-specific reasoning traces, Reinforcement Learning (RL) with execution accuracy and novel text2SQL rewards, and a combined SFT+RL approach—across multiple benchmark datasets.
Our findings answer three research questions. RQ1: while general-purpose reasoning in pretrained LLMs offers limited benefits for complex Text2SQL under ZSL, incorporating task-specific reasoning traces via SFT significantly improves performance, particularly for smaller models. RQ2: RL proved highly effective across all models and datasets, especially for queries demanding multi-hop reasoning - pure RL often yielded the best performance on challenging subsets. RQ3: the introduced dense rewards for RL training, based on the QATCH metrics, outperformed traditional sparse rewards (EX), enhancing the model’s ability to learn from complex reasoning tasks. RQ4: the combined SFT+RL strategy demonstrated strong generalization across diverse datasets, suggesting it strikes an effective balance between learning general reasoning patterns (via SFT) and optimizing for task-specific correctness (via RL). Our Think2SQL-7B model, trained with RL and the introduces dense reward $R_{\text{QATCH}}$, achieved performance surpassing models with more than 400 billion parameters on the BIRD dataset, showcasing the power of targeted reasoning reinforcement.
Despite these promising results, this study has limitations. We deliberately isolated the SQL generation process by providing the relevant database schema subset, excluding the challenge of automated schema linking, which is critical in real-world applications. Furthermore, the SFT training relied on synthetically generated reasoning traces, whose style and quality might influence outcomes. The core experiments focused on the Qwen-Coder-2.5 family, and findings might vary across different model architectures.
Future work could explore the impact of combining sparse and dense rewards in reinforcement learning, aiming to better balance signal strength and specificity during training. Investigating the effectiveness of these reasoning-centered training strategies across a broader range of LLM architectures and model scales, as well as on diverse datasets—including proprietary or domain-specific corpora—would further validate their generalizability. Additionally, a more in-depth qualitative analysis of model failure cases could yield valuable insights into how different reward strategies influence performance on complex or nuanced queries.